answers are shown,. please show how to solve this type of problem with ti 83

ai The pulse rates of 156 randomly selected adult males vary from a low of 41 bpm to a high of 113 bpm. Find the minimum sample size required to estimate the mean pulse rate of adult males. Assume that we want 90% confidence that the sample mean is within 2 bpm of the population mean. Complete parts (a)through(c) below. a. Find the sample size using the range rule of thumb to estimate ơ n 220 Round up to the nearest whole number as needed.) b. Assume that ơ-10.3 bpm, based on the value s-10.3 bpm from the sample of 156 male pulse rates. n 72 Round up to the nearest whole number as needed.) c. Compare the results from parts (a) and (b). Which result is likely to be better The result from part (a) is the result from part (b). The result from Vis likely to be better because 0
Answer #1

SolutionA:

range rule of thumb to estimate standard deviation

=max-min/4

=113-41/4

=72/4

=18

n=18

MARKED ANSWER WRONG

SolutionB:

n=(Z*sigma/E)^2

=(1.645*10.3/2)^2

=71.77

=72

n=72

MARKED ANSWER IS CORRECT

Solutionc:

the results from part A is different from part B.the result from part B is better because uses margin of error E